### v2

# Deep Partially Linear Transformation Model for Right-Censored Survival Data

Abstract
Although the Cox proportional hazards model is well established and extensively used in the analysis of survival data, the proportional hazards (PH) assumption may not always hold in practical scenarios. The semiparametric transformation model extends the conventional Cox model and also includes many other survival models as special cases. This paper introduces a deep partially linear transformation model (DPLTM) as a general and flexible framework for estimation, inference and prediction. The proposed method is capable of avoiding the curse of dimensionality while still retaining the interpretability of some covariates of interest. We derive the overall convergence rate of the maximum likelihood estimators, the minimax lower bound of the nonparametric deep neural network (DNN) estimator, the asymptotic normality and the semiparametric efficiency of the parametric estimator. Comprehensive simulation studies demonstrate the impressive performance of the proposed estimation procedure in terms of both estimation accuracy and prediction power, which is further validated by an application to a real-world dataset.
Keywords: deep learning; minimax lower bound; monotone B-splines; partially linear transformation model; semiparametric efficiency.
1 Introduction
The Cox proportional hazards model (Cox, 1972, 1975) is probably the most widely used semiparametric model in survival analysis. However, the Cox model assumes proportional hazards for individuals, which may be too simplistic and often violated in practice. An example is the acquired immune deficiency syndrome (AIDS) data assembled by the U.S. Center for Disease Control, which includes 295 blood transfusion patients diagnosed with AIDS prior to July 1, 1986. One primary interest is to explore the effect of age at transfusion on the induction time, but Grigoletto and Akritas (1999) revealed that the PH assumption fails with such data even with the use of the reverse time PH model. The class of semiparametric transformation models emerges as a more general and flexible alternative that requires no prior assumption and has recently received tremendous attention. Most of the frequently employed survival models can be viewed as specific cases of transformation models, including the Cox proportional hazards model, the proportional odds model (Bennett, 1983), the accelerated failure time model (Wei, 1992) and the usual Box-Cox model. Multiple estimation procedures have been thoroughly discussed for transformation models with right-censored data (Cheng et al., 1995; Chen et al., 2002), current status data (Zhang et al., 2013), interval-censored data (Zeng et al., 2016), competing risk data (Fine, 1999) and recurrent event data (Zeng and Lin, 2007).
Linear transformation models allow interpretation for all covariate effects, but one limitation is that the linearity assumption is sometimes too unrealistic for complex real-world data. For instance, in the New York University Women’s Health Study (NYUWHS), a question of our interest is whether the time of developing breast carcinoma is influenced by the sex hormone levels, and a strongly nonlinear relationship between them is identified by Zeleniuch-Jacquotte et al. (2004). To accommodate linear and nonlinear covariate effects simultaneously, partially linear transformation models were developed (Ma and Kosorok, 2005; Lu and Zhang, 2010) and later generalized to the case with varying coefficients (Li et al., 2019; Al-Mosawi and Lu, 2022). Nevertheless, these works either only consider the simple case of univariate nonlinear effect, or assume the nonparametric effects to be additive, which is often inconsistent with the reality.
Public health and clinical studies in the age of big data have benefited substantially from large-scale biomedical research resources such as UK Biobank and the Surveillance, Epidemiology, and End Results (SEER) program. Such databases often contain dozens of or more covariates of interest to be handled simultaneously. Much important information would be left out if data from these sources is fitted by the simple linear or partially linear additive model. Recently, deep learning has rapidly evolved into a dominant and promising method in a wide range of sectors involving high-dimensional data such as computer vision (Krizhevsky et al., 2012), natural language processing (Collobert et al., 2011) and finance (Heaton et al., 2017). Deep neural networks have also brought about significant advancements in survival analysis. They have been combined with a variety of survival models like the Cox proportional hazards model (Katzman et al., 2018; Zhong et al., 2022), the cause-specific model for competing risk data (Lee et al., 2018), the cure rate model (Xie and Yu, 2021a, b) and the accelerated failure time model (Norman et al., 2024).
Statistical theory of deep learning associates its empirical success with its strong capability to approximate functions from specific spaces (Yarotsky, 2017; Schmidt-Hieber, 2020). Inspired by this, Zhong et al. (2022) considered DNNs for estimation in a partially linear Cox model, and developed a general theoretical framework to study the asymptotic properties of the partial likelihood estimators. This pioneering work has been extended to the cases of current status data (Wu et al., 2024) and interval-censored data (Du et al., 2024). In this paper, we propose a deep partially linear transformation model for highly complicated right-censored survival data. Some covariates of our primary interest are modelled linearly to keep their interpretability, while other covariate effects are approached by a deep ReLU network to alleviate the curse of dimensionality. The convergence rate of the estimators given by maximizing the log likelihood function is free of the nonparametric covariate dimension under proper conditions and faster than those using traditional methods like kernels or splines. Additionally, the parametric and nonparametric estimators are proved to be semiparametric efficient and minimax rate-optimal respectively.
The rest of the paper is organized as follows. In Section 2, we introduce the framework of our proposed method and the maximum likelihood estimation procedure based on deep neural networks and monotone splines. Section 3 is devoted to establishing the asymptotic properties of the estimators. In Section 4, we conduct extensive simulation experiments to examine the finite sample performance of the proposed model and make comparisons with other models. An application to a real-world dataset is provided in Section 5. Section 6 concludes the paper. Detailed proofs of lemmas and theorems, along with further numerical results are given in the Supplementary Materials.
2 Methodology
2.1 Likelihood function
We consider a study of $n$ subjects with right-censored survival data, where the survival time and the censoring time are denoted by $U$ and $C$, respectively. $\bm{Z}$ is a $p$-dimensional covariate vector impacting on the survival time linearly, and $\bm{X}$ is a $d$-dimensional covariate vector whose effects will be modelled nonparametrically. In the presence of censoring, the observations consist of $n$ i.i.d copies $\{\bm{V}_{i}=(T_{i},\Delta_{i},\bm{Z}_{i},\bm{X}_{i}),\ i=1,\cdots,n\}$ from $\bm{V}=(T,\Delta,\bm{Z},\bm{X})$, where $T=\min\left\{U,C\right\}$ is the observed event time and $\Delta=I(U\leq C)$ is the censoring indicator, with $I(\cdot)$ being the indicator function. It is generally assumed in survival analysis that $U$ is independent of $C$ conditional on $(\bm{Z},\bm{X})$.
To model the effects of the covariates $(\bm{Z},\bm{X})\in\mathbb{R}^{p}\times\mathbb{R}^{d}$ on the survival time $U$, the partially linear transformation model specifies that
$H(U)=-\bm{\beta}^{\top}\bm{Z}-g(\bm{X})+\varepsilon,$ (1)
where $H$ is an unknown transformation function assumed to be strictly increasing, $\bm{\beta}\in\mathbb{R}^{p}$ denotes the unspecified parametric coefficients and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is an unknown nonparametric function. To simplify our notation, we denote the parameters to be estimated by $\bm{\eta}=(\bm{\beta},H,g)$, and assume that the joint distribution of $(\Delta,\bm{Z},\bm{X})$ is free of $\bm{\eta}$. $\varepsilon$ is an error term with a known continuous distribution function independent of $(\bm{Z},\bm{X})$.
Many useful survival models are included in the class of partially linear transformation models as special cases. For example, (1) reduces to the partially linear Cox model or the partially linear proportional odds model when $\varepsilon$ follows the extreme value distribution or the standard logistic distribution, respectively. If we choose $H(t)=\log t$, (1) serves as the partially linear accelerated failure time (AFT) model. When $\varepsilon$ follows the normal distribution and there is no censoring, (1) generalizes the partially linear Box-Cox model.
Let ($f_{\varepsilon}$, $S_{\varepsilon}$, $\lambda_{\varepsilon}$, $\Lambda_{\varepsilon}$) and ($f_{U}$, $S_{U}$, $\lambda_{U}$, $\Lambda_{U}$) be the probability density function, survival function, hazard function and cumulative hazard function of $\varepsilon$ and $U$, respectively. Then it is straightforward to verify that
$f_{U}(t|\bm{Z},\bm{X})=H^{\prime}(t)f_{\varepsilon}(H(t)+\bm{\beta}^{\top}\bm{Z}+g(\bm{X})),\ S_{U}(t|\bm{Z},\bm{X})=S_{\varepsilon}(H(t)+\bm{\beta}^{\top}\bm{Z}+g(\bm{X})),$
$\lambda_{U}(t|\bm{Z},\bm{X})=H^{\prime}(t)\lambda_{\varepsilon}(H(t)+\bm{\beta}^{\top}\bm{Z}+g(\bm{X})),\ \Lambda_{U}(t|\bm{Z},\bm{X})=\Lambda_{\varepsilon}(H(t)+\bm{\beta}^{\top}\bm{Z}+g(\bm{X})).$
Therefore, the observed information of a single object under (1) can be expressed as
$\mathcal{L}(\bm{V})$ =$=\left\{f_{U}(T|\bm{Z},\bm{X})\right\}^{\Delta}\left\{S_{U}(T|\bm{Z},\bm{X})\right\}^{1-\Delta}q(\Delta,\bm{Z},\bm{X})$
=$=\left\{\lambda_{U}(T|\bm{Z},\bm{X})\right\}^{\Delta}\exp\left\{-\Lambda_{U}(T|\bm{Z},\bm{X})\right\}q(\Delta,\bm{Z},\bm{X})$
=$=\left\{H^{\prime}(T)\lambda_{\varepsilon}(H(T)+\bm{\beta}^{\top}\bm{Z}+g(\bm{X}))\right\}^{\Delta}\exp\left\{-\Lambda_{\varepsilon}(H(T)+\bm{\beta}^{\top}\bm{Z}+g(\bm{X}))\right\}q(\Delta,\bm{Z},\bm{X}),$
where $q(\Delta,\bm{X},\bm{Z})$ is the joint density of $(\Delta,\bm{X},\bm{Z})$. Then the log likelihood function of $\bm{\eta}=(\bm{\beta},H,g)$ given $\{\bm{V}_{i}=(T_{i},\Delta_{i},\bm{Z}_{i},\bm{X}_{i}),\ i=1,\cdots,n\}$ can be written as
$L_{n}(\bm{\eta})=\frac{1}{n}\sum_{i=1}^{n}\Big{\{}\Delta_{i}\log H^{\prime}(T_{i})+\Delta_{i}\log\lambda_{\varepsilon}(H(T_{i})+$ $\bm{\beta}^{\top}\bm{Z}_{i}+g(\bm{X}_{i}))$ (2)
$-\Lambda_{\varepsilon}(H(T_{i})+\bm{\beta}^{\top}\bm{Z}_{i}+g(\bm{X}_{i}))\Big{\}}.$
2.2 Sieve Maximum Likelihood Estimation
To achieve a faster convergence rate of the estimators, two different function spaces of growing capacity with respect to the sample size $n$ for the infinite-dimensional parameters $g$ and $H$ are chosen for the estimation procedure.
For the estimation of the nonparametric function $g$, we use a sparse deep ReLU network space with depth $K$, width vector $\bm{p}=(p_{0},\cdots,p_{K+1})$, sparsity constraint $s$ and norm constraint $D$, which has been specified in Schmidt-Hieber (2020) and Zhong et al. (2022) as
$\mathcal{G}(K,\bm{p},s,D)=$$\Bigg{\{}g(x)=(W_{K}\sigma(\cdot)+v_{K})\circ\cdots\circ(W_{0}x+v_{0}):\mathbb{R}^{p_{0}}\mapsto\mathbb{R}^{p_{K+1}},\Big{.}$
.$\quad\Big{.}W_{k}\in\mathbb{R}^{p_{k+1}\times p_{k}},\ v_{k}\in\mathbb{R}^{p_{k+1}},\ \max\left\{\left\lVert W_{k}\right\rVert_{\infty},\left\lVert v_{k}\right\rVert_{\infty}\right\}\leq 1\text{ for }k=0,\cdots,K,\Big{.}$
.$\quad\Big{.}\sum_{k=0}^{K}\left(\left\lVert W_{k}\right\rVert_{0}+\left\lVert v_{k}\right\rVert_{0}\right)\leq s,\ \left\lVert g\right\rVert_{\infty}\leq D\Bigg{\}},$
where $W_{k}$ and $v_{k}$ are the weight and bias of the $(k+1)$-th layer of the network, respectively, $\sigma(x)=\max\left\{x,0\right\}$ is the ReLU activation function operating component-wise on a vector, $\left\lVert\cdot\right\rVert_{0}$ denotes the number of non-zero entries of a vector or matrix, and $\left\lVert\cdot\right\rVert_{\infty}$ denotes the sup-norm of a vector, matrix or function.
To estimate the strictly increasing transformation function $H$, a monotone spline space is adopted. We assume that the support of the observed event time $T$ is in a closed interval $[L_{T},U_{T}]$ with $0<L_{T}<U_{T}<\tau$, where $\tau$ is the end time of the study, and partition the interval $[L_{T},U_{T}]$ into $K_{n}+1$ sub-intervals with respect to the knot set
$\Upsilon=\left\{L_{T}=t_{0}<t_{1}<\cdots<t_{K_{n}+1}=U_{T}\right\},$
then we can construct $q_{n}=K_{n}+\ell+1$ B-spline basis functions $B_{k}(t|\ell),\ k=1,\cdots,q_{n}$ that are piecewise polynomials of degree $\ell$ and span the space of polynomial splines $\mathcal{S}$ of degree $\ell$ with $\Upsilon$. We usually set $K_{n}=O(n^{\nu})$ and $\underset{1\leq k\leq K_{n}+1}{\max}|t_{k}-t_{k-1}|=O(n^{-\nu})$ for some $20<\nu<1/$ 2 based on theoretical analysis, and $\ell\geq 2$ so that the spline function is at least continuously differentiable. Besides, by Theorem 5.9 of Schumaker (2007), it suffices to impose the monotone increasing condition on the coefficients of basis functions to ensure the monotonicity of the spline function. Thus we consider the following function space $\Psi$ which is a subset of $\mathcal{S}$:
$\Psi=\left\{\sum_{j=1}^{q_{n}}\gamma_{j}B_{j}(t|\ell):-\infty<\gamma_{1}\leq\cdots\leq\gamma_{q_{n}}<\infty,\ t\in[L_{T},U_{T}]\right\}.$
We denote the true value of $\bm{\eta}=(\bm{\bm{\beta}},H,g)$ by $\bm{\eta}_{0}=(\bm{\bm{\beta}}_{0},H_{0},g_{0})$, then $\bm{\eta}_{0}$ is estimated by maximizing the log likelihood function (2):
$\widehat{\bm{\eta}}=(\widehat{\bm{\bm{\beta}}},\widehat{H},\widehat{g})=\underset{(\bm{\bm{\beta}},H,g)\in\mathbb{R}^{p}\times\Psi\times\mathcal{G}}{\operatorname*{arg\,max}}L_{n}(\bm{\bm{\beta}},H,g),$ (3)
where $\mathcal{G}=\mathcal{G}(K,\bm{p},s,\infty)$. However, it may be challenging to perform gradient-based optimization algorithms with the monotoncity-constrained coefficients. We consider using a reparameterizaion approach with $\widetilde{\gamma}_{1}=\gamma_{1}$ and $\widetilde{\gamma}_{j}=\log(\gamma_{j}-\gamma_{j-1})$ for $2\leq j\leq q_{n}$ to enforce monotonicity, and then conduct optimization with respect to $\left\{\widetilde{\gamma}_{j}\right\}_{j=1}^{q_{n}}$ instead.
3 Asymptotic Properties
In this section, we describe the asymptotic properties of the log likelihood estimators in (3) under appropriate conditions. First, we impose some restrictions on the true nonparametric function $g_{0}$. Recall that a Hölder class of smooth functions with parameters $\alpha$, $M$ and domain $\mathbb{D}\subset\mathbb{R}^{d}$ is defined as
$\mathcal{H}_{d}^{\alpha}(\mathbb{D},M)=\left\{g:\mathbb{D}\mapsto\mathbb{R}:\sum_{\kappa:|\kappa|<\alpha}\left\lVert\partial^{\kappa}g\right\rVert_{\infty}+\sum_{\kappa:|\kappa|=\lfloor\alpha\rfloor}\sup_{x,y\in\mathbb{D},x\neq y}\frac{|\partial^{\kappa}g(x)-\partial^{\kappa}g(y)|}{\left\lVert x-y\right\rVert_{\infty}^{\alpha-\lfloor\alpha\rfloor}}\leq M\right\}.$
We further consider a composite smoothness function space that has been introduced in Schmidt-Hieber (2020):
$\mathcal{H}(q,\bm{\alpha},\bm{d},\widetilde{\bm{d}},M):=\Big{\{}g=$ $g_{q}\circ\cdots\circ g_{0}:g_{i}=(g_{i1},\cdots,g_{id_{i+1}})^{\top}\text{ and }$
$g_{ij}\in\mathcal{H}^{\alpha_{i}}_{\widetilde{d}_{i}}([a_{i},b_{i}]^{\widetilde{d}_{i}},M)\text{, for some }|a_{i}|,|b_{i}|<M\Big{\}}.$
Furthermore, we denote $\widetilde{\alpha}_{i}=\alpha_{i}\prod_{k=i+1}^{q}(\alpha_{k}\wedge 1)$ and $\delta_{n}=\max_{i=0,\cdots,q}n^{-\widetilde{\alpha}_{i}/(2\widetilde{\alpha}_{i}+\widetilde{d}_{i})}$, and the following regularity assumptions are required to derive asymptotic properties:
(C1) $K=O(\log n)$, $s=O(n\delta_{n}^{2}\log n)$ and $n\delta_{n}^{2}\lesssim\min(p_{k})_{k=1,\cdots,K}\leq\max(p_{k})_{k=1,\cdots,K}\lesssim n$.
(C2) The covariates $(\bm{Z},\bm{X})$ take value in a bounded subset of $\mathbb{R}^{p+d}$ with joint probability density function bounded away from zero. Without loss of generality, we assume that the domain of $\bm{X}$ is $[0,1]^{d}$. Moreover, the parameter $\bm{\bm{\beta}}_{0}$ lies in a compact subset of $\mathbb{R}^{p}$.
(C3) The nonparametric function $g_{0}$ is in $\mathcal{H}_{0}=\{g\in\mathcal{H}(q,\bm{\alpha},\bm{d},\widetilde{\bm{d}},M):\mathbb{E}\{g(\bm{X})\}=0\}$.
(C4) The $k$-th derivative of the transformation function $H_{0}$ is Lipschitz continuous for any $k\geq 1$. Particularly, its first derivative is strictly positive on $[L_{T},U_{T}]$.
(C5) The hazard function of the error term $\lambda_{\varepsilon}$ is log-concave and twice continuously differentiable on $\mathbb{R}$. Besides, its first derivative is strictly positive on compact intervals.
(C6) There is some constant $\xi>0$ such that $\mathbb{P}(\Delta=1|\bm{Z},\bm{X})>\xi$ and $\mathbb{P}(U\geq\tau|\bm{Z},\bm{X})>\xi$ almost surely with respect to the probability measure of $(\bm{Z},\bm{X})$.
(C7) The sub-density $p(t,\bm{x},\Delta=1)$ of $(T,\bm{X},\Delta=1)$ is bounded away from zero and infinity on $[0,\tau]\times[0,1]^{d}$.
(C8) For some $k>1$, the $k$-th partial derivative of the sub-density $p(t,\bm{x},\bm{z},\Delta=1)$ of $(T,\bm{X},\bm{Z},\Delta=1)$ with respect to $(t,\bm{x})$ exists and is bounded on $[0,\tau]\times[0,1]^{d}$.
Condition (C1) configures the structure of the function space $\mathcal{G}(K,\bm{p},s,D)$ by specifying its hyperparameters which grow with the sample size. Condition (C2) is commonly used for semiparametric partially linear models. Condition (C3) yields the model’s identifiability. Technical conditions (C4)-(C6) are utilized to establish the consistency and the convergence rate of estimators. It is worth noting that the seemingly strong assumptions in Condition (C5) are satisfied by many familiar survival models such as the Cox proportional hazards model, the proportional odds model and the Box-Cox model. Condition (C7) guarantees the existence of the information bound for $\bm{\bm{\beta}}_{0}$. Condition (C8) establishes the asymptotic normality of $\widehat{\bm{\bm{\beta}}}$.
For any $\bm{\eta}_{1}=(\bm{\bm{\beta}}_{1},H_{1},g_{1})$ and $\bm{\eta}_{2}=(\bm{\bm{\beta}}_{2},H_{2},g_{2})$, define
$d(\bm{\eta}_{1},\bm{\eta}_{2})=\left\{\|\bm{\bm{\beta}}_{1}-\bm{\bm{\beta}}_{2}\|^{2}+\|g_{1}-g_{2}\|^{2}_{L^{2}([0,1]^{d})}+\|H_{1}-H_{2}\|^{2}_{L^{2}([0,\tau])}+\|H^{\prime}_{1}-H^{\prime}_{2}\|^{2}_{L^{2}([0,\tau])}\right\}^{1/2},$
where $\|\bm{\bm{\beta}}_{1}-\bm{\bm{\beta}}_{2}\|^{2}=\sum_{i=1}^{p}(\beta_{i1}-\beta_{i2})^{2}$, $\|g_{1}-g_{2}\|^{2}_{L^{2}([0,1]^{d})}=\mathbb{E}\left\{g_{1}(\bm{X})-g_{2}(\bm{X})\right\}^{2}$, $\|H_{1}-H_{2}\|^{2}_{L^{2}([0,\tau])}=\mathbb{E}\left\{H_{1}(T)-H_{2}(T)\right\}^{2}$ and $\|H^{\prime}_{1}-H^{\prime}_{2}\|^{2}_{L^{2}([0,\tau])}=\mathbb{E}\left[\Delta\left\{H^{\prime}_{1}(T)-H^{\prime}_{2}(T)\right\}^{2}\right]$. For $\bm{\eta}=(\bm{\bm{\beta}},H,g)$ and $\bm{V}=(T,\Delta,\bm{Z},\bm{X})$, define $\phi_{\bm{\eta}}(\bm{V})=H(T)+\bm{\bm{\beta}}^{\top}\bm{Z}+g(\bm{X})$ and $\Phi_{\bm{\eta}}(\bm{V})=\Delta\frac{\lambda_{\varepsilon}^{\prime}(\phi_{\bm{\eta}}(\bm{V}))}{\lambda_{\varepsilon}(\phi_{\bm{\eta}}(\bm{V}))}-\lambda_{\varepsilon}(\phi_{\bm{\eta}}(\bm{V})).$ Then we have the following theorems whose proofs are provided in the Supplementary Materials:
Theorem 1 (Consistency and rate of convergence).
Suppose conditions (C1)–(C6) hold, and it holds that $(2w+1)^{-1}<\nu<(2w)^{-1}$ for some $w\geq 1$, then
$d(\widehat{\bm{\eta}},\bm{\eta}_{0})=O_{p}(\delta_{n}\log^{2}n+n^{-w\nu}).$
Therefore, the proposed DNN-based model is able to mitigate the curse of dimensionality and enjoys a faster rate of convergence than traditional nonparametric methods such as kernels or splines when the intrinsic dimension $\widetilde{\bm{d}}$ is relatively low.
Furthermore, the minimax lower bound for the estimation of $g_{0}$ is presented below:
Theorem 2 (Minimax lower bound).
Suppose conditions (C1)-(C6) hold. Define $\mathbb{R}^{p}_{M}=\left\{\bm{\bm{\beta}}\in\mathbb{R}^{p}:\|\bm{\bm{\beta}}\|\leq M\right\}$, then there exists a constant $0<c<\infty$, such that
$\underset{\widehat{g}}{\inf}\underset{(\bm{\bm{\beta}}_{0},H_{0},g_{0})\in\mathbb{R}^{p}_{M}\times\Psi\times\mathcal{H}_{0}}{\sup}\mathbb{E}\left\{\widehat{g}_{0}(\bm{X})-g_{0}(\bm{X})\right\}^{2}\geq c\delta_{n}^{2},$
where the infimum is taken over all possible estimators $\widehat{g}$ based on the observed data.
Next theorem gives the efficient score and information bound for $\bm{\bm{\beta}}_{0}$. To start with, let $\Psi_{H_{0}}$ denote the collection of all subfamilies $\left\{H_{s_{1}}\in L^{2}([0,\tau]:s_{1}\in(-1,1)\right\}$ such that $\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(H_{s_{1}}-H_{0})-a_{1}\|_{L^{2}([0,\tau])}=0$ and $\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(\log H_{s_{1}}^{\prime}-\log H_{0}^{\prime})-a_{2}\|_{L^{2}([0,\tau])}=0$, where $a_{1},a_{2}\in L^{2}([0,\tau])$, and define
$\mathbb{T}_{H_{0}}=\Big{\{}a_{1}\in L^{2}([0,\tau]):\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(H_{s_{1}}-H_{0})-a_{1}\|_{L^{2}([0,\tau])}=0$ for some subfamily
$\left\{H_{s_{1}}:s_{1}\in(-1,1)\right\}\in\Psi_{H_{0}}\Big{\}},$
$\mathbb{T}_{H^{\prime}_{0}}=\Big{\{}a_{2}\in L^{2}([0,\tau]):\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(\log H_{s_{1}}^{\prime}-\log H_{0}^{\prime})-a_{2}\|_{L^{2}([0,\tau])}$ $=0\text{ for some subfamily }$
$\left\{H_{s_{1}}:s_{1}\in(-1,1)\right\}\in\Psi_{H_{0}}\Big{\}}.$
Similarly, let $\mathcal{H}_{g_{0}}$ denote the collection of all subfamilies $\left\{g_{s_{2}}\in L^{2}([0,1]^{d}):s_{2}\in(-1,1)\right\}\subset\mathcal{H}_{0}$ such that $\lim_{s_{2}\rightarrow 0}\|s_{2}^{-1}(g_{s_{2}}-g_{0})-b\|_{L^{2}([0,1]^{d})}=0$ with $b\in L^{2}([0,1]^{d})$, and define
$\mathbb{T}_{g_{0}}=\Big{\{}b\in L^{2}([0,1]^{d}):\lim_{s_{2}\rightarrow 0}\|s_{2}^{-1}(g_{s_{2}}-g_{0})-b\|_{L^{2}([0,1]^{d})}=0$ for some subfamily
$\left\{g_{s_{2}}:s_{2}\in(-1,1)\right\}\in\mathcal{H}_{g_{0}}\Big{\}}.$
Let $\overline{\mathbb{T}}_{H_{0}}$, $\overline{\mathbb{T}}_{H^{\prime}_{0}}$ and $\overline{\mathbb{T}}_{g_{0}}$ be the closed linear span of $\mathbb{T}_{H_{0}}$, $\mathbb{T}_{H^{\prime}_{0}}$ and $\mathbb{T}_{g_{0}}$, respectively.
Theorem 3 (Efficient score and information bound).
Suppose conditions (C2)-(C7) hold, the efficient score for $\bm{\bm{\beta}}_{0}$ is
$\ell^{*}_{\bm{\bm{\beta}}}(\bm{V};\bm{\eta}_{0})=[\bm{Z}-\bm{a}_{1*}(T)-\bm{b}_{*}(\bm{X})]\Phi_{\bm{\eta}_{0}}(\bm{V})-\Delta\bm{a}_{2*}(T),$
where $(\bm{a}_{1*}^{\top},\bm{a}_{2*}^{\top},\bm{b}_{*}^{\top})^{\top}\in\overline{\mathbb{T}}_{H_{0}}^{p}\times\overline{\mathbb{T}}_{H^{\prime}_{0}}^{p}\times\overline{\mathbb{T}}_{g_{0}}^{p}$ is the least favorable direction minimizing
$\mathbb{E}\left\{\left\|\left\{\bm{Z}-\bm{a}_{1}(T)-\bm{b}(\bm{X})\right\}\Phi_{\bm{\eta}_{0}}(\bm{V})-\Delta\bm{a}_{2}(T)\right\|^{2}_{c}\right\},$
with the notation $\|\cdot\|^{2}_{c}$ denoting the component-wise square of a vector. Moreover, the information bound for $\bm{\bm{\beta}}_{0}$ is
$I(\bm{\beta}_{0})=\mathbb{E}\left\{\ell^{*}_{\bm{\bm{\beta}}}(\bm{V};\bm{\eta}_{0})\right\}^{\bigotimes 2}.$
The last theorem states that, though the overall convergence rate is slower than $n^{-1/2}$, we can still derive the asymptotic normality of $\widehat{\bm{\bm{\beta}}}$ with $\sqrt{n}$-consistency.
Theorem 4 (Asymptotic Normality).
Suppose conditions (C1)-(C8) hold. If $(2w+1)^{-1}<\nu<(2w)^{-1}$ for some $w\geq 1$, $I(\bm{\beta}_{0})$ is nonsingular and $n\delta_{n}^{4}\rightarrow 0$, then
[…]
4 Simulation Studies
We conduct simulation studies in this section to investigate the finite sample performance of the proposed DPLTM method, and compare it with the linear transformation model (LTM) and the partially linear additive transformation model (PLATM) with respect to both estimation and prediction. The DPLTM method is implemented by PyTorch (Paszke et al., 2019).
We first provide some computational details. The models are fitted by maximizing the log likelihood function (2) with respect to the parameters $\bm{\beta}$, $\widetilde{\gamma}_{j}$’s, $W_{k}$’s and $v_{k}$’s. For the DPLTM method, all the parameters are contained in one framework and simultaneously updated in each epoch, where the Adam optimizer (Kingma and Ba, 2014) is employed due to its efficiency and reliability. All components of $\bm{\beta}$ and all $\widetilde{\gamma}_{j}$’s are initialized to 0 and -1, respectively, while PyTorch’s default initialization algorithm is utilized for $W_{k}$’s and $v_{k}$’s.
The hyper-parameters in the simulations, including the number of hidden layers, the number of neurons in each hidden layer, the number of epochs, learning rate (Goodfellow, 2016), dropout rate (Srivastava et al., 2014) and the number of B-spline basis functions are tuned based on the log likelihood on the validation data via a grid search. We set the number of neurons in each hidden layer to be identical for convenience. We use cubic splines (i.e. $\ell$=3) to estimate $H_{0}(t)$ to achieve sufficient smoothness, with the number of interior knots $K_{n}$ chosen in the ranges $\lfloor n^{1/3}\rfloor,\cdots,2\lfloor n^{1/3}\rfloor$, and the number of basis functions $q_{n}=K_{n}+\ell+1$ can be determined. Candidates for other hyperparameters are presented in the Supplementary Materials.
To avoid overfitting, we use the strategy of early stopping (Goodfellow, 2016) in the training process. Specifically, if the log likelihood on the validation data stops increasing for a predetermined number of consecutive epochs, which is an indication of overfitting, we then terminate the training and obtain the estimates.
In all simulations, the linearly modelled covariates $\bm{Z}$ have two independent components, where the first is generated from a Bernoulli distribution with the success probability 0.5, and the second follows a normal distribution with both mean and variance 0.5. The covariate vector with nonlinear effects $\bm{X}$ are 5-dimensional and from a Gaussian copula with correlation coefficient 0.5. Each coordinate of $\bm{X}$ is assumed to be uniformly distributed on $[0,2]$. We take the true treatment effect $\bm{\beta}_{0}=(1,-1)$ and consider the following three designs for the true nonparametric function $g_{0}(\bm{x})$ with $\bm{x}\in[0,2]^{5}$:
  • Case 1 (Linear): $g_{0}(\bm{x})=0.25(x_{1}+2x_{2}+3x_{3}+4x_{4}+5x_{5}-15)$,
  • Case 2 (Additive): $g_{0}(\bm{x})=1.85\big{\{}\sin(2\pi x_{1})+\cos(\pi x_{2}/2)/2+\log(x_{3}^{2}+1)/3+(x_{4}-x_{4}^{3})/4+(e^{x_{5}}-1)/5-0.428\big{\}}$,
  • Case 3 (Deep): $g_{0}(\bm{x})=2.2\big{\{}\sin(2\pi x_{1}x_{2})+\cos(\pi x_{2}x_{3}/2)/2+\log(x_{3}x_{4}+1)/3+(x_{4}-x_{3}x_{4}x_{5})/4+(e^{x_{5}}-1)/5-0.778\big{\}}$.
The three cases correspond to LTM, PLATM and DPLTM respectively. The intercept terms -15, -0.428 and -0.778 guarantee that $\mathbb{E}g_{0}(\bm{X})=0$ in all cases, and the factors 0.25, 1.85 and 2.2 are used to scale the signal ratio $\text{Var}\left\{g_{0}(\bm{X})\right\}/\text{Var}\left\{\bm{\beta}_{0}^{\top}\bm{Z}\right\}$ within $[5,7]$.
The distribution of the error term $\varepsilon$ is chosen from the class of logarithm transformations (Dabrowska and Doksum, 1988), whose hazard function is of the form
$\lambda(t)=\frac{e^{t}}{1+re^{t}},$
with $r=0,0.5,1$. Actually, $r=0$ and $r=1$ correspond to the proportional hazards model and the proportional odds model respectively. Note that all three candidates satisfy the condition (C5) in our theoretical analysis.
The true transformation function $H_{0}(t)$ is set respectively to be $\log t$ for $r=0$, $\log(2e^{0.5t}-2)$ for $r=0.5$ and $\log(e^{t}-1)$ for $r=1$. Then we can generate survival time $U$ via its distribution function $F_{U}(t)=F_{\varepsilon}(H_{0}(t)+\bm{\beta}_{0}^{\top}\bm{Z}+g_{0}(\bm{X}))$ based on the inverse transform method. The censoring time is generated from a uniform distribution on $(0,c_{0})$, where the constant $c_{0}$ is chosen to approximately achieve the pre-specified censoring rate of 40% and 60% ($c_{0}=$2.95 or 0.85 for $r=0$, $c_{0}=$2.75 or 0.9 for $r=0.5$, $c_{0}=$2.55 or 1 for $r=1$, all kept the same for the three different cases of the underlying function $g_{0}(\bm{x})$).
We conduct 200 runs of simulation in each scenario with the sample size $n$=1000 and 2000. Our observations consist of $\{\bm{V}_{i}=(T_{i},\Delta_{i},\bm{Z}_{i},\bm{X}_{i}),\ i=1,\cdots,n\}$, where $T_{i}=\min\left\{U_{i},C_{i}\right\}$ and $\Delta_{i}=I(U_{i}\leq C_{i})$. We randomly split the samples into training data (80%) and validation data (20%), the former is used to fit models and obtain estimates, and the latter is utilized to tune the hyper-parameters. In addition, We generated $n_{\text{test}}=$200 and 400 test samples (corresponding to the cases of $n=$1000 and 2000 respectively) that are independent of the training samples for evaluation.
To estimate the asymptotic covariance matrix $I(\bm{\beta}_{0})^{-1}$ for inference, we estimate the least favorable direction $(\bm{a}_{1*},\bm{a}_{2*},\bm{b}_{*})$ by minimizing the empirical version of the objective function given in Theorem 3:
$(\widehat{\bm{a}}_{1*},\widehat{\bm{a}}_{2*},\widehat{\bm{b}}_{*})=\underset{(\bm{a}_{1},\bm{a}_{2},\bm{b})}{\arg\min}\frac{1}{n}\sum_{i=1}^{n}\left\|[\bm{Z}_{i}-\bm{a}_{1}(T_{i})-\bm{b}(\bm{X}_{i})]\Phi_{\widehat{\bm{\eta}}}(\bm{V}_{i})-\Delta_{i}\bm{a}_{2}(T_{i})\right\|^{2}_{c}.$
We approximate $(\bm{a}_{1*},\bm{a}_{2*},\bm{b}_{*})$ using another DNN due to the absence of closed-form expressions for them. The information bound $I(\bm{\beta}_{0})$ can then be estimated by
$\widehat{I}(\bm{\beta}_{0})=\frac{1}{n}\sum_{i=1}^{n}\Big{\{}[\bm{Z}_{i}-\widehat{\bm{a}}_{1*}(T_{i})-\widehat{\bm{b}}_{*}(\bm{X}_{i})]\Phi_{\widehat{\bm{\eta}}}(\bm{V}_{i})-\Delta_{i}\widehat{\bm{a}}_{2*}(T_{i})\Big{\}}^{\bigotimes 2}.$
For evaluation of the performance of $\widehat{g}$, we compute the relative error (RE) based on the test data, which is given by
$\text{RE}(\widehat{g})=\left\{\frac{\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\left[\left\{\widehat{g}(\bm{X}_{i})-\overline{\widehat{g}}\right\}-g_{0}(\bm{X}_{i})\right]^{2}}{\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\{g_{0}(\bm{X}_{i})\}^{2}}\right\}^{1/2},$
where $\overline{\widehat{g}}=\sum_{i=1}^{n_{\text{test}}}\widehat{g}(\bm{X}_{i})/n_{\text{test}}$.
We also evaluate and compare the predictive ability of the three methods using the concordance index (C-index) by Harrell et al. (1982). C-index is one of the most commonly used metrics to assess the predictive power of models in survival analysis. It measures the probability that the predicted survival times preserve the ranks of true survival times, which is defined as
$\text{C}=\mathbb{P}(\widehat{T}_{i}<\widehat{T}_{j}|T_{i}<T_{j},\Delta_{i}=1),$
where $\widehat{T}_{i}$ denotes the predicted survival time of the $i$-th individual. Larger C-index values indicate better predictive performance. For the semiparametric transformation model, the C-index can be empirically calculated as
$\widehat{\text{C}}=\frac{\sum_{i=1}^{n_{\text{test}}}\sum_{j=1}^{n_{\text{test}}}\Delta_{i}1(T_{i}\leq T_{j})1(\widehat{\bm{\beta}}\bm{Z}_{i}+\widehat{g}(\bm{X}_{i})\geq\widehat{\bm{\beta}}\bm{Z}_{j}+\widehat{g}(\bm{X}_{j}))}{\sum_{i=1}^{n_{\text{test}}}\sum_{j=1}^{n_{\text{test}}}\Delta_{i}1(T_{i}\leq T_{j})}.$
The bias and standard deviation of the parametric estimate $\widehat{\bm{\beta}}$ derived from 200 simulation runs are presented in Table 1. It can be seen that the proposed DPLTM method provides asymptotically unbiased estimates in all scenarios. The biases for DPLTM are sometimes slightly higher than those for LTM and PLATM under Case 1, and PLATM under Case 2 respectively, which is expected because these two cases perfectly satisfy the simple linear and additive settings respectively. However, DPLTM greatly outperforms LTM and PLATM under Case 3 with a highly complicated true nonparametric function $g_{0}$, where the other two models are remarkably more biased than DPLTM and do not improve with increasing sample size. Moreover, the empirical standard deviation decreases steadily as $n$ increases for all three models under each simulation setting.
Tables 2 shows the empirical coverage probabilities of 95% confidence intervals built with the asymptotic variance of $\widehat{\bm{\beta}}$ derived from the estimated information bound $\widehat{I}(\bm{\beta}_{0})$. It is clear that the coverage rate of DPLTM is generally close to the nominal level of 95%, while PLATM gives inferior results under Case 3 and LTM shows poor coverage under both Case 2 and Case 3 because of the large bias.
Table 3 reports the relative error averaged over 200 simulation runs and its standard deviation of the norparametric estimates $\widehat{g}$ on the test data. Likewise, the DPLTM estimator shows consistently strong performance in all three cases, and the metric gets smaller as the sample size increases. In contrast, LTM and PLATM behave poorly when the underlying function does not coincide with their respective model assumptions, which implies that they are unable to provide accurate estimates of complicated nonparametric functions.
Table 4 exhibits the average and standard deviation of C-index on the test data based on 200 simulation runs. Unsurprisingly, predictions obtained by the DPLTM method are comparable to or only a little worse than those by LTM and PLATM in simple settings, but DPLCM shows great superiority over the other two models under the complex Case 3 as it produces much more precise estimates for both $\bm{\beta}$ and $g$ in this scenario.
In the Supplementary Materials, we further measure the accuracy in estimating the transformation function $\widehat{H}$ with the weighted integrated squared error (WISE) and assess the calibration of the three methods using the integrated calibration index (ICI) (Austin et al., 2020). Similar conclusions can be drawn from results on these metrics. More importantly, we provide a comprehensive comparison between our method with the DPLCM method introduced by Zhong et al. (2022), and demonstrate that our method is generally more powerful and reliable than DPLCM due to its flexibility. In conclusion, these simulation results manifest that our DPLTM method performs consistently well with respect to both estimation and prediction in various settings.
5 Application
In this section, we apply the proposed DPLTM method to real-world data to demonstrate its outstanding performance. We employ lung cancer data from the Surveillance, Epidemiology, and End Results (SEER) database. The SEER database is a publicly available cancer reporting system supported by the National Cancer Institute, and enables the study of trends in cancer incidence, prevalence and survival in the United States. We select patients who were diagnosed with lung cancer in 2015, with age between 18 and 85 years old, survival time longer than one month and received treatment no more than 730 days (2 years) after diagnosis. We extract 5 discrete covariates that serve as $\bm{Z}$, including gender, marital status, primary cancer, separate tumor nodules in ipsilateral lung, chemotherapy, and 5 continuous covariates that are treated as $\bm{X}$, including age, time from diagnosis to treatment in days, CS tumor size, CS extension and CS lymph nodes. Samples with missing covariates are discarded, which results in a dataset consisting of 28950 subjects with a censoring rate of 25.63%. The dataset is split into a training set, a validation set and a test set with a ratio of 64:16:20. The main purpose of our study is to assess the predictive performance of our DPLTM method while still allowing interpretation for some covariates.
The candidates for the error term distribution of transformation models are the same as in simulation studies, i.e. the class of logarithm transformations with $r=0,0.5,1$. To obtain better predictions, we have to select the ‘optimal’ model. We calculate the log likelihood on the validation data under the three fitted models for the DPLTM method, which are -6618.40, -6469.49 and -6440.13 for $r$=0, 0.5 and 1 respectively. This suggest that the model with $r=1$, i.e. the partially linear proportional odds model fits this dataset best and is then used for interpretation and prediction.
The estimated coefficients (EST) of the linear component, along with their estimated standard errors (ESE) and corresponding asymptotic $p$-values for the DPLTM method with $r=1$ are given in Table 12. It is easy to see that all linearly modelled covariates except the one indicating whether it is a primary cancer are statistically significant. To be specific, females, the married, patients without separate tumor nodules in ipsilateral lung and those who received chemotherapy after diagnosis have significantly longer survival times.
As in simulation studies, the data is fitted by LTM, PLATM and DPLTM, all corresponding to the case of $r=1$, to compare their predictive power, and we use C-index as the evaluation criterion. The C-index for the DPLTM method are 0.7028, which is notably larger than that for LTM (0.6466) and PLATM (0.6675) respectively. This shows the considerably superior predictive validity of DPLTM due to its ability to deal with highly complicated survival data in practical scenarios.
In the Supplementary Materials, we also assess the prediction accuracy of the three models with the integrated calibration index (ICI), and further compare DPLTM with the DPLCM method in terms of prediction using both C-index and ICI on this dataset. To summarize, these results exhibit that our model is more robust on real-world data as well.
6 Conclusion
This paper introduces a DPLTM method for right-censored survival data. It combines deep neural networks with partially linear transformation models, which encompass a number of useful models as specific cases. Our method demonstrates outstanding predictive performance while maintaining interpretability of the parametric component. The maximum likelihood estimators converge at a rate that depends only on the intrinsic dimension, and we establish the semiparametric efficiency of coefficient estimators and the minimax lower bound of the deep neural network estimator respectively. Numerical results show that DPLTM not only significantly outperforms the simple linear and additive models, but also offers major improvements over the DPLCM method.
This paper has only focused on semiparametric transformation models for right-censored survival data. It is straightforward to extend our methodology to other survival models like the cure rate model (Kuk and Chen, 1992; Lu and Ying, 2004), and other types of survival data such as current status data and interval-censored data. Moreover, multimodal data, such as gene sequences and histopathological images, have brought new insights into survival analysis. It is thus of great importance to combine our method with more advanced neural network architectures like deep convolutional neural networks (LeCun et al., 1989), deep residual networks (He et al., 2016) and transformer networks (Vaswani, 2017) for unstructured data, and develop a more general theoretical framework for them.